Mat Aneli SC 133 Spring9-MULLER Activities and Due Dates ) Resources Check Answer Consider a cloudless day on which the sun shines down across the United States. If 2542 J of energy reaches a square meter m2) of the United States in one hour, how much total solar energy reaches the entire United States per hour? The entire area of the United States is 9,158960 km kWh MacBook Pro 838 Ul
Answer #1

Total surface area of United States=9158960 km2 =9158960*106 m2

2542 KJ of energy reaches per m2 area in one hour

thus total energy reached in one hour= (9158960*106)m2*(2542)KJ m-2 h-1

= 23,282,076,320*106 KJ h-1

=23,282,076,320,000,000KJ h-1